# LLMs for Literature Review: Are we there yet?

### v1

@@ 3.4 Retrieval and Re-ranking Experiments
We use an ensemble of search engines to retrieve candidates based on an abstract. We now describe the search engines used in our approach. Based on Table 1, we select the Semantic Scholar (S2) API as the search engine to retrieve search results using LLM-generated keywords. As discussed in Section 3.2, we also explore SPECTER2 embeddings for retrieval and compare it with the different prompting-based strategies.
@@
$\text{Normalized Recall}=\frac{|\text{Retrieved}\cap\text{Ground Truth}|}{|\text{Ground Truth}|};\text{Precision@k}=\frac{|\text{Retrieved}\cap\text{Ground Truth}|}{k}$ (1)
@@
where $k$ is the number of top-$k$ recommendations, “Retrieved” denotes the set of top-$k$ candidate papers, and “Ground Truth” denotes the set of Ground Truth (GT) cited papers (papers cited by the query paper).
@@
In Table 2, we examine the behaviour of the GPT-4 reranking approach in more detail. Using GPT-4 as a reranker in the manner discussed above, produces an incomplete list 41% of the time. In 3% cases, GPT-4 produces a list which has repeated values with some rare cases with garbage values or numbers (e.g. 2020). We conclude that this strategy of using GPT-4 for reranking is brittle.
@@ 3.5 Positive Effect of Attribution Verification
We conduct an ablation study on the first 100 papers of the larger set of 500, and focus on the top $k=40$ papers, which is representative of the typical number of papers cited in the Machine Learning community. In Figure 3 we show the result of removing the verification step in our debate re-ranking strategy, i.e. in this ablation we do not check if the sentences extracted by the LLM are indeed present in the candidate paper abstract. We find that removing this kind of attribution verification leads to a drop in the precision and normalized recalls, especially for lower $k$ values. We perform a t-test to test the significance of the drop and find that for both precision and normalized recall, the drop is significant, with $p$-values of $44.7e-$ 4 and $61.9e-$ 6 for precision and normalized recall curves, respectively. This indicates that proper attribution also allows the LLM to provide more accurate ranking of candidates.
Figure 3: The effect of removing the referenced content verification step in our debate ranking strategy. We plot precision and normalized recall for two variants of the debate ranking strategy. For this ablation study, we select a smaller subset of $n=100$ query abstracts, set $k=40$, and repeat the experiment for three random seeds. We plot the mean and show the standard deviation as the shaded region. We find that the precision and normalized recall drop slightly upon removing the verification step. This different is significant (as determined by t-test,) indicating that the verification step is crucial for the success of the debate ranking strategy.
@@ 4 Literature Review Generation
4.1 Plan Based Generation Approach & Model Variants
We now focus on generating the related work section of a scientific document from a user-supplied list of papers. In a real-world scenario, this list might be obtained through traditional means, from the above-mentioned automated methods, or some combination. We evaluate several dimensions of writing quality in the generated text. Importantly, while modern LLMs can yield seemingly well-written text passages, “hallucinations” remain a problem and can be particularly egregious when LLMs are used in scientific writing (Athaluri et al., 2023). The hallucination of statements not entailed by the contents of cited papers and the hallucinations of imaginary papers that do not exist is a well-known issue of even the most powerful LLMs. We use ideas from retrieval augmented generation (RAG) techniques (Lewis et al., 2020) and instruction prompting to address the key problem of hallucinations.Our work also aims to increase the number of papers from the desired set that are indeed discussed (the coverage).
We present our general framework and problem setup in Figure 4. We use the abstract of a query paper — the one for which we generate a literature review, along with the abstracts of the set of papers to be cited (the retrieved abstracts of reference papers) to generate the related work section of the query paper. Our approach relies on prompting LLMs in different ways and measuring the similarity of the generated literature review text to ground truth literature reviews found within a corpus of recent scientific papers — i.e. ones not used in the training set of the underlying LLMs. We use both automated methods and human evaluations in our experiments below.
We propose to further decompose the writing task to increase passage quality, factual accuracy, and coverage. We examine different strategies for generating a writing plan, a line-by-line description including citations of the passage to write. These writing plans also give authors (users) control over the output passages. This is likely essential in practice to meet author preferences and possible publication constraints. These plans are defined and generated in such a way that the user can interact with and edit them if desired, or they can be used as an automatically generated intermediate (but human understandable) representation. We now we describe our proposed methods, elements of their use in practice and their evaluation in more detail.
@@ 4.1 Plan Based Generation Approach & Model Variants
In what we shall refer to as plan-based generation (Plan), a model is prompted with a known (user provided) plan to produce X sentences in Y words and cite references on specific lines. These plans are obtained from the ground truth data and serve as a proxy for a user who desires text to be written with specific constraints. This kind of Plan strongly guides generated text to what the user desires (as observed by the final ground truth text). During evaluation this might be thought of as a form of teacher-forcing at the structural level. An example of the format of these plans is provided below:
@@
Prompted plan The previous method replicates the scenario of a detailed user provided plan. However, one can also generate such plans automatically from an LLM. The model is prompted to first generate a plan of sentences and citations which it would then condition upon to generate the final related work text. When used as an interactive tool we envision the user might start with a suggested plan, see the corresponding generated full literature review text, and then iteratively edit the plan and regenerate the result. See our Appendix (Figures 9, 10 and 11) for the different prompts used in our experiments. We also experiment with 2 other strategies in which researchers could prompt the model:
Per cite We first use a two-stage strategy to generate content relevant to each cited paper. In the first stage, the model is prompted to generate related works in 1-2 lines for each individual reference citation. All the outputs for different citations are combined together to form the generated related work. In the second stage, the LLM summarizes and paraphrases the output of the first stage.
Sentence by sentence Based on the Ground truth (GT) related work and the citation on each line, we prompt the model to generate one sentence conditioned on the abstract, the reference cited in that line, and the generated draft so far. In the absence of a citation or at the start, the model is prompted only with abstract and draft of the generated work till now.
5 Generation Experiments
For the following studies on generating related work, we introduce an additional corpus. We extend the Multi-XScience corpus (Lu et al., 2020) to include the full text of research papers. We also reuse the RollingEval-Aug introduced in Section 3. We use HuggingFace Transformers (Wolf et al., 2019) and PyTorch (Paszke et al., 2017) for our experiments [^7] Code will be released at github.com and calculate ROUGE scores (Lin, 2004) using the Huggingface evaluate library. Details on the dataset and the implementation are in Appendix B and D, respectively. Similar to Lu et al. (2020), we extract the ground truth cited references as the relevant papers and evaluate only the generated outputs from different systems.
@@ 5.1 Generation Baselines
Extractive baselines As in Lu et al. (2020), we report the performance of LexRank (Erkan & Radev, 2004) and TextRank (Mihalcea & Tarau, 2004). We also create a simple one-line extractive baseline which extracts the first line of the abstract and combines all the citations together to form the output.
Abstractive finetuned baselines We use the model outputs of Hiersum (Liu & Lapata, 2019) and Pointer-Generator (See et al., 2017) from Lu et al. (2020) for abstractive finetuned baselines. We also reproduce the finetuned PRIMER (Xiao et al., 2021) model (considered to be SOTA).
@@
Open and closed source models We use different chat versions (7B, 13B, 70B) of Llama 2-Chat [^8] We refer to Llama 2-Chat models as Llama 2 and GPT-3.5-turbo as GPT-3.5 for brevity. (Touvron et al., 2023) as zero-shot open-source LLM baselines. For closed-source models, we evaluate zero-shot both GPT-3.5-turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023). Since they perform best in our initial evaluations, we use the closed-source models in combination with the different generation strategies (Per-cite, Sentence by sentence, plan-based, and learned plan) from Section 4.1.
@@ 6 Results and Observations
From Table 3, we first note that unsupervised extractive models provide a strong baseline compared to abstractive 0-shot single document summarization baselines. Fine-tuning these abstractive models on Multi-XScience (released initially with the benchmark) improves performance at least to the level of extractive models. We reproduce the PRIMER model using their open-source code but find results lower than reported. As such, we consider the Pointer-Generator method to be the current state-of-the-art (SOTA).
@@
CodeLlama 34B-Instruct 22.608 5.032 12.553
CodeLlama 34B-Instruct (Plan) 27.369 5.829 14.705
@@
Llama 2-Chat 13B 23.998 5.472 12.923
Llama 2-Chat 70B 23.769 5.619 12.745
@@
GPT-3.5-turbo (Plan) 30.192 7.028 15.551
GPT-4 (Plan) 33.044 7.352 17.624
@@
Single-document summarizers (LongT5, Flan T5) perform poorly in the zero-shot settings with limited ability to cite references. We are limited in the prompt we can provide (because of the training prompts) and resort to ‘Summarize the text and cite sources.’ Galactica’s performance is encouraging compared to other models in the same group, but inspecting its output reveals that it generates the whole introduction of the paper instead of the related work. The model is very sensitive to the prompts used (mostly as suffixes) and struggles to follow instructions. Falcon 180-B, on the other hand, tends to hallucinate user turns and considers this task as multiple turns of user-system exchange, even though we prompted to generate relevant outputs.
@@
Our teacher-forced plan-based framework improves the scores over the 0-shot baseline for both closed-sourced (GPT-3.5 and GPT-4) and open-sourced LLMs with Llama 2 70B achieving similar scores as GPT-3.5 on both the original Multi-XScience and the new RollingEval-Aug dataset (in Table 4). Llama 2 70B gets more uplift with the plan compared to GPT models where manual inspection reveals fewer hallucinations in the outputs (see qualitative results in Table 12 in Appendix using our abstract). In Table 5, we evaluate controllability of LLM based generation using sentence plans and find that GPT-4 tends to follow the plan more closely. It follows the exact plan 60% of the time, often producing fewer sentences than provided. Llama 2 70B comes in second place in following the plan instructions and GPT-3.5 struggles to follow the plan precisely. We also experiment with a learned plan strategy where the model first generates a plan and then autoregressively generates the output. Though it improves the results over 0-shot baseline, it does not outperform the teacher-forced plan generation in terms of automatic metrics. There is a considerable drop in performance on RollingEval-Aug dataset compared to the original Multi-XScience in terms of ROUGE1/2. It gives more credibility to the hypothesis that the Multi-XScience test set is in the training data of these LLMs and/or that there is a shift in the distribution of these more recent papers. Nevertheless, we find similar scores and ranking patterns between models as for Multi-XScience. We provide other experiments related to fine-tuning, longer context and CodeLLM in the Appendix C.
@@
Figure 5: Human evaluation study where annotators ranked the generations of 0-shot models with their sentence-plan based counterparts. On the Y-axis, we show counts from an overall sample size of 58 annotations for Llama 2-Chat and 54 for GPT-4 (where ranking ties are allowed). We see a reduction of 58.6% cases of hallucinations to 32.7% for Llama 2-Chat and 29.6% to 11.6% for GPT-4 using plan based prompting.
We also run a human evaluation study using 6 expert annotators. They rank model generated outputs for 160 papers with 3 citations [^10] This was done to reduce cognitive load on annotators to read abstracts of 4+ papers and 2 model outputs. Annotators have research experience in machine learning at the Masters or Doctorate level. where we show the abstract of the query paper, cited references, and the model outputs of 0-shot vs plan counterparts side-by-side. We randomly selected 80 examples each for both GPT-4 and Llama 2-Chat comparison. Experts had the flexibility to choose rank 1 for one or both the models, otherwise they could select rank 2 for each of them. We also ask the annotators to identify hallucinations i.e. which model generated content not from the abstracts. The interface is described further and shown in the Appendix (Figure 8). Out of 160, we find agreement among the annotators for 112 examples (54 for GPT4 and 58 for Llama 2-Chat). The results in Figure 5 show that humans rank generated response to be significantly better [^11] We used McNemar test (Lachenbruch, 2014) to measure statistical significance. for Llama 2-Chat Plan based model where it was ranked at the top 32 times compared to 18 times for 0-shot. We can also observe a similar trend for GPT-4, where annotators rank GPT-4 Plan-based output 37 times at Rank 1 compared to 30 for GPT-4 0-shot. In terms of hallucinations, we find significant reductions in the hallucination for plan-based Llama 2-Chat models from 34 to 19 instances, where the 0-shot model often provided a made-up citation (XYZ et al.), possibly from background knowledge. Only 6 cases of hallucination were found for GPT-4 plan compared to 16 instances for 0-shot vanilla GPT-4 model.

### v2

@@ 3.4 Retrieval and Re-ranking Experiments
We use an ensemble of search engines to retrieve candidates based on an abstract. We now describe the search engines used in our approach. Based on Table 1, we select the Semantic Scholar (S2) API as the search engine to retrieve search results using LLM-generated keywords. While the SERP API provides access to a broader set of search results, it is expensive, making it less practical for large-scale retrieval. Additionally, the S2 API is specifically designed for academic literature, offering strong performance comparable to the SERP API. As discussed in Section 3.2, we also explore SPECTER2 embeddings for retrieval and compare it with the different prompting-based strategies.
@@
[…]
@@
where “Retrieved” denotes the set of all candidate papers, “Retrieved@k” denotes the set of top $k$ candidate papers, and “Ground Truth” denotes the set of papers cited by the query paper. Unlike standard recall, which measures how many ground truth citations are retrieved, Normalized Recall@k measures how effectively the retrieval method prioritizes the most relevant papers in the top-$k$. By normalizing over the total relevant papers retrieved at any rank, this metric helps evaluate ranking quality independently of retrieval coverage. We include a working example in Appendix D.1 to further demonstrate the difference between normalized recall and standard recall.
@@
In Table 2, we examine the behaviour of the GPT-4 reranking approach in more detail. Using GPT-4 as a reranker in the manner discussed above produces an incomplete list 41% of the time. In 3% cases, GPT-4 produces a list with repeated values and, in some rare cases, with garbage values or numbers (e.g. 2020). We conclude that this strategy of using GPT-4 for reranking is brittle.
@@ 3.5 Positive Effect of Attribution Verification
[…]
Figure 3: The effect of removing the referenced content verification step in our debate ranking strategy. We plot precision and normalized recall for two variants of the debate ranking strategy. For this ablation study, we select a smaller subset of $n=100$ query abstracts, set $k=40$, and repeat the experiment for three random seeds. We plot the mean and show the standard deviation as the shaded region. We find that the precision and normalized recall drop slightly upon removing the verification step. This difference is significant (as determined by the t-test,) indicating that the verification step is crucial for the success of the debate ranking strategy.
@@ 4 Literature Review Generation
Plan Based Generation Approach & Model Variants. We now focus on generating the related work section of a scientific document from a user-supplied list of papers. In a real-world scenario, this list might be obtained through traditional means, from the above-mentioned automated methods, or some combination. We evaluate several dimensions of writing quality in the generated text. Importantly, while modern LLMs can yield seemingly well-written text passages, “hallucinations” remain a problem and can be particularly egregious when LLMs are used in scientific writing (Athaluri et al., 2023). The hallucination of statements not entailed by the contents of cited papers and the hallucinations of imaginary papers that do not exist is a well-known issue of even the most powerful LLMs. We use ideas from retrieval augmented generation (RAG) techniques (Lewis et al., 2020) and instruction prompting to address the key problem of hallucinations.Our work also aims to increase the number of papers from the desired set that are indeed discussed (the coverage).
We present our general framework and problem setup in Figure 4. We use the abstract of a query paper — the one for which we generate a literature review, along with the abstracts of the set of papers to be cited (the retrieved abstracts of reference papers) to generate the related work section of the query paper. For evaluation, our approach relies on prompting LLMs in different ways and measuring the similarity of the generated literature review text to ground truth literature reviews found within a corpus of recent scientific papers — i.e. ones not used in the training set of the underlying LLMs. We use both automated methods and human evaluations in our experiments below.
We propose to further decompose the writing task to increase passage quality, factual accuracy, and coverage. We examine different strategies for generating a writing plan, a line-by-line description including citations of the passage to write. These writing plans also give authors (users) control over the output passages. This is likely essential in practice to meet author preferences and possible publication constraints. These plans are defined and generated in such a way that the user can interact with and edit them if desired, or they can be used as an automatically generated intermediate (but human-understandable) representation. We now describe our proposed methods, their use in practice and their evaluation in more detail.
@@
In what we refer to as plan-based generation (Plan), a model is prompted with a known (user-provided) plan to produce X sentences in Y words and cite references on specific lines. These plans are obtained from the ground truth data and serve as a proxy for a user who desires text to be written with specific constraints. This kind of Plan strongly guides generated text to the user’s desires (as observed by the final ground truth text). During evaluation, this might be considered a form of teacher-forcing at the structural level. An example of the format of these plans is provided below:
@@
Prompted plan. The previous method replicates the scenario of a detailed user-provided plan. However, one can also generate such plans automatically from an LLM. The model is prompted first to generate a plan of sentences and citations, which it would then condition upon to generate the final related work text. When used as an interactive tool, we envision the user might start with a suggested plan, see the corresponding generated full literature review text, and then iteratively edit the plan and regenerate the result. See our Appendix (Figures 9, 10 and 11) for the differences in the prompts used in our experiments. We also experiment with two other strategies in which researchers could prompt the model:
Per cite. We first use a two-stage strategy to generate content relevant to each cited paper. In the first stage, the model is prompted to generate related works in 1-2 lines for each individual reference citation. All the outputs for different citations are combined together to form the generated related work. In the second stage, the LLM summarizes and paraphrases the output of the first stage.
Sentence by sentence Based on the Ground truth (GT) related work and the citation on each line, we prompt the model to generate one sentence conditioned on the abstract, the reference cited in that line, and the generated draft so far. In the absence of a citation or at the start, the model is prompted only with the abstract and draft of the generated work till now.
5 Generation Experiments.
For the following studies on generating related work, we introduce an additional corpus. We extend the Multi-XScience corpus (Lu et al., 2020) to include the full text of research papers. We also reuse the RollingEval-Aug introduced in Section 3. We use HuggingFace Transformers (Wolf et al., 2019) and PyTorch (Paszke et al., 2017) for our experiments [^7] Code can be accessed at https://github.com/LitLLM/litllms-for-literature-review-tmlr and calculate ROUGE scores (Lin, 2004) using the Huggingface’s evaluate library. Details on the dataset and the implementation are in Appendix B and D, respectively. Similar to Lu et al. (2020), we extract the ground truth cited references as the relevant papers and evaluate only the generated outputs from different systems. Since ROUGE score only measures token-level similarity and does not account for semantic meaning, we also report BERTScore and Llama-3-Eval in Table 4. We use BERTScore, an embedding-based evaluation metric, to account for the semantic meaning of two texts during evaluation as well. On the other hand, we use Llama-3-Eval, an open-source variant of the widely used G-Eval metric (Liu et al., 2023b), as G-Eval has been shown to correlate better with human preferences.
@@ 5.1 Generation Baselines
Extractive baselines As in Lu et al. (2020), we report the performance of LexRank (Erkan & Radev, 2004) and TextRank (Mihalcea & Tarau, 2004). We also create a simple one-line extractive baseline which extracts the first line of the abstract and combines all the citations to form the output.
Abstractive finetuned baselines We use the model outputs of Hiersum (Liu & Lapata, 2019) and Pointer-Generator (See et al., 2017) from Lu et al. (2020) for abstractive finetuned baselines. We also reproduce the finetuned PRIMER (Xiao et al., 2021) model (considered to be the SOTA).
@@
Open and closed source models We use different chat versions (7B, 13B, 70B) of Llama 2-Chat [^8] We refer to Llama 2-Chat models as Llama 2 and GPT-3.5-turbo as GPT-3.5 for brevity. (Touvron et al., 2023) as zero-shot open-source LLM baselines. For closed-source models, we evaluate zero-shot both GPT-3.5-turbo (Brown et al., 2020) and GPT-4 (OpenAI, 2023). Since they perform best in our initial evaluations, we use the closed-source models in combination with the different generation strategies (Per-cite, Sentence by sentence, plan-based, and learned plan) from Section 4.
@@ 6 Results and Observations
From Table 3, we first note that unsupervised extractive models provide a strong baseline compared to abstractive 0-shot single document summarization baselines. Fine-tuning these abstractive models on Multi-XScience (released initially with the benchmark) improves performance at least to the level of extractive models. We reproduce the PRIMER model using their open-source code but find lower-than-reported results. As such, we consider the Pointer-Generator method to be the current state-of-the-art (SOTA).
@@
CodeLlama 34B-Instruct 22.608 5.032 12.553 82.418 66.898
CodeLlama 34B-Instruct (Plan) 27.369 5.829 14.705 83.386 67.362
@@
Llama 2-Chat 13B 23.998 5.472 12.923 82.855 69.237
Llama 2-Chat 70B 23.769 5.619 12.745 82.943 70.980
@@
GPT-3.5-turbo (Plan) 30.192 7.028 15.551 84.203 72.729
GPT-4 (Plan) 33.044 7.352 17.624 85.151 75.240
@@
Single-document summarizers (LongT5, Flan T5) perform poorly in the zero-shot settings with limited ability to cite references. We are limited in the prompt we can provide (because of the training prompts) and resort to “Summarize the text and cite sources.” Galactica’s performance is encouraging compared to other models in the same group, but inspecting its output reveals that it generates the whole introduction of the paper instead of the related work. The model is very sensitive to the prompts used (mostly as suffixes) and struggles to follow instructions. Falcon 180-B, on the other hand, tends to hallucinate user turns and considers this task as multiple turns of user-system exchange, even though we prompted to generate relevant outputs.
@@
Our teacher-forced plan-based framework improves the scores over the 0-shot baseline for both closed-sourced (GPT-3.5 and GPT-4) and open-sourced LLMs, with Llama 2 70B achieving similar scores as GPT-3.5 on both the original Multi-XScience and the new RollingEval-Aug dataset (in Table 4). In Table 4, we also notice that BERTScore and Llama-3-Eval exhibit the same trends as ROUGE scores except in the case where GPT-3.5-turbo (Plan) obtains a higher Llama-3-Eval score than Llama 2-Chat 70B (Plan). These values also showcase the weaker discerning power of BERTScore compared to Llama-3-Eval as all the models achieve a high BERTScore between 82-85%. Llama 2 70B gets more uplift with the plan compared to GPT models where manual inspection reveals fewer hallucinations in the outputs (see qualitative results in Table 13 in Appendix using our abstract). In Table 5, we evaluate the controllability of LLM-based generation using sentence plans and find that GPT-4 tends to follow the plan more closely. It follows the exact plan 60% of the time, often producing fewer sentences than provided. Llama 2 70B comes in second place in following the plan instructions and GPT-3.5 struggles to follow the plan precisely. We also experiment with a learned plan strategy where the model first generates a plan and then autoregressively generates the output. Though it improves the results over 0-shot baseline, it does not outperform the teacher-forced plan generation in terms of automatic metrics. There is a considerable drop in performance on RollingEval-Aug dataset compared to the original Multi-XScience in terms of ROUGE1/2. It gives more credibility to the hypothesis that the Multi-XScience test set is in the training data of these LLMs and/or that there is a shift in the distribution of these more recent papers. Nevertheless, we find similar scores and ranking patterns between models as for Multi-XScience. We provide other experiments related to fine-tuning, longer context, and CodeLLM in Appendix C, and we provide cost estimates for different methods in Appendix D.5.
@@
Figure 5: Human evaluation study where annotators ranked the generations of 0-shot models with their sentence-plan-based counterparts. On the Y-axis, we show counts from an overall sample size of 58 annotations for Llama 2-Chat and 54 for GPT-4 (where ranking ties are allowed). We see a reduction of 58.6% cases of hallucinations to 32.7% for Llama 2-Chat and 29.6% to 11.6% for GPT-4 using plan-based prompting.
We also run a human evaluation study using 6 expert annotators. They rank model generated outputs for 160 papers with 3 citations [^10] This was done to reduce cognitive load on annotators to read abstracts of 4+ papers and 2 model outputs. Annotators have research experience in machine learning at the Masters or Doctorate level. where we show the abstract of the query paper, cited references, and the model outputs of 0-shot vs plan counterparts side-by-side. We randomly selected 80 examples each for GPT-4 and Llama 2-Chat comparisons. Experts had the flexibility to choose rank 1 for one or both the models, otherwise they could select rank 2 for each. We also ask the annotators to identify hallucinations, i.e. which model generated content not from the abstracts. The interface is described further and shown in the Appendix (Figure 8). Out of 160, we find agreement among the annotators for 112 examples (54 for GPT4 and 58 for Llama 2-Chat). The results in Figure 5 show that humans rank generated response to be significantly better [^11] We used McNemar test (Lachenbruch, 2014) to measure statistical significance. for Llama 2-Chat Plan based model where it was ranked at the top 32 times compared to 18 times for 0-shot. We can also observe a similar trend for GPT-4, where annotators rank GPT-4 Plan-based output 37 times at Rank 1 compared to 30 for GPT-4 0-shot. In terms of hallucinations, we find significant reductions in the hallucination for plan-based Llama 2-Chat models from 34 to 19 instances, where the 0-shot model often provided a made-up citation (XYZ et al.), possibly from background knowledge. Only 6 cases of hallucination were found for GPT-4 plan compared to 16 instances for 0-shot vanilla GPT-4 model.